# How to explain grokking

### v2

@@ 1.1 SGLD and Eyring’s formula
The Eyring formula of kinetic theory describes the reaction rate (transition between two potential wells due to diffusion of the type (2)): the reaction rate is proportional to
@@
High entropy of the potential well lowers the free energy, that is, wide potential wells better capture the particle.
@@ 2 Grokking and thermodynamics
Grokking (delayed generalization) phenomenon in learning was discovered in [8], see also [9], [10], [11]. Overparameterized network was trained to perform modular arithmetics (in particular, addition and other operations modulo 97). The phenomenon of delayed generalization was found, and the following observations were made:
@@
The zero-risk manifold contains narrows (with low entropy) and wide valleys (with high entropy). The correct solution for learning problem (”structure”, for example algorithm of modular addition) mostly likely lies in the high-entropy region of the zero-risk manifold.
@@
Grokking is a random walk in the zero risk manifold $f(\{z\},x)=0$ (or Brownian motion). For Brownian motion the path traveled is proportional to the root of the number of steps tsimilar-$\sim\sqrt{t}$. By the second law of thermodynamics the system will travel to areas of high entropy at the zero risk manifold.
@@
Grokking is the transition from the initial region (part of the zero-risk manifold where the Brownian motion of grokking starts) to the grokking region (a valley with high entropy, a neighborhood of the solution of the learning problem in the form of a ”structure”).
The Eyring’s formula $e^{-\beta(F_{1}-F_{0})}$ gives an approximation of the dependence of the reciprocal of the grokking time on the training sample size. Here $F_{0}$ and $F_{1}$ are free energies of the initial region for grokking and of the transition state (the neighborhood of the solution). The entropy of the initial region decreases linearly with the training sample size (hence the free energy $F_{0}=E_{0}-\theta S_{0}$ increases linearly). The free energy $F_{1}$ of the transition region changes slightly with the training sample size. Consequently, the grokking time increases exponentially when the training sample decreases (this agrees with observations in [8]).
@@ 3 Summary
We also have explained the grokking (delayed generalization) phenomenon and some properties of grokking observed in [8].

### v3

@@ Abstract
1 Introduction
Stochastic gradient descent, which is a basic model in machine learning, when taken in the form of the stochastic gradient Langevin dynamics, is a direct analogue of a random walk along the reaction coordinates, which is a principal model in chemical kinetics. It is natural to use in learning the ideas of kinetics and thermodynamics, in particular, if the gradient descent is an analogue of the energy driven chemical reaction, then the stochastic gradient descent should include also the discussion of entropy, i.e. of the free energy minimization. We show that the transition to generalization in grokking is an analogue of the entropy driven chemical reaction (see the discussion below).
@@ 2.1 SGLD and Eyring’s formula
Remark. Let us discuss a short introduction to statistical mechanics for mathematicians, based on arguments from [3]. Let us put (2) as
${\partial u\over\partial t}=-\theta e^{-\beta f/2}A^{*}Ae^{\beta f/2}u,\quad A=e^{-\beta f/2}\nabla e^{\beta f/2}.$
The operator $A^{*}A$ is positive, moreover
$e^{-t\theta e^{-\beta f/2}A^{*}Ae^{\beta f/2}}=e^{-\beta f/2}e^{-t\theta A^{*}A}e^{\beta f/2},$
$e^{-t\theta A^{*}A}$ is a contraction semigroup in $L_{2}$. The condition $e^{\beta f/2}u\in L_{2}$ is
[…]
For the stationary solution $e^{-\beta f}$ eq. (3) is the integrability of the Gibbs distribution
[…]
zero is the eigenvalue of $A^{*}A$
$\left[e^{-\beta f/2}A^{*}Ae^{\beta f/2}\right]e^{-\beta f}=0.$
Then if the initial condition $u_{0}$ satisfies (3) and Gibbs distribution is integrable (4), the dynamics $e^{-t\theta A^{*}A}e^{\beta f/2}u_{0}$ belongs to $L_{2}$ and converges to $e^{-\beta f/2}$, and the solution $e^{-\beta f/2}e^{-t\theta A^{*}A}e^{\beta f/2}u_{0}$ of (2) converges to Gibbs distribution $e^{-\beta f}$.
The Eyring formula of kinetic theory describes the reaction rate (for the transition between two potential wells due to diffusion of the type (2)): the reaction rate is proportional to
@@
High entropy of the potential well lowers the free energy, that is, wide potential wells better capture the particle for the dynamics (2).
@@ 2.2 SGLD and learning
Therefore, we consider the stochastic gradient optimization as a minimization of free energy of a potential well in statistical mechanics, or an analogue of chemical reaction.
Gradient descent optimization is a chemical reaction by energy minimization, but in general also entropic contribution to free energy is important.
@@ 3 Grokking and thermodynamics
Grokking (delayed generalization) phenomenon in learning was discovered in [9], see also [10], [11], [12], [13]. Overparameterized network was trained to perform modular arithmetics (in particular, addition and other operations modulo 97). The phenomenon of delayed generalization was found, and the following observations were made [9]:
@@
The zero-risk manifold contains narrows, or ravines (areas with low entropy) and wide valleys (with high entropy). For ravine, or river–valleys landscape, learning is a motion along the ravine (the reaction coordinate in chemistry). The correct solution for learning problem (”structure”, for example algorithm of modular addition) mostly likely lies in the high-entropy region of the zero-risk manifold. The ravine method [15] was developed to improve optimization for such a landscape.
@@
Grokking is a random walk in the zero risk manifold $f(\{z\},x)=0$ (or Brownian motion). For Brownian motion the path traveled is proportional to the root of the number of steps tsimilar-$\sim\sqrt{t}$.
By the second law of thermodynamics the system will travel to areas of high entropy at the zero risk manifold, i.e. the transition to generalization in grokking is an analogue of the entropic mechanism of chemical reaction (moreover the energy difference for different points in the zero risk manifold equals to zero).
@@
Grokking is the transition from the initial region (part of the zero-risk manifold where the Brownian motion of grokking starts) to the grokking region (a valley with high entropy, a neighborhood of the solution of the learning problem in the form of a ”structure”) along the ”reaction coordinate” in the zero-risk manifold.
The Eyring’s formula $e^{-\beta(F_{1}-F_{0})}$ gives an approximation of the dependence of the reciprocal of the grokking time on the training sample size. Here $F_{0}$ and $F_{1}$ are free energies of the initial region for grokking and of the transition state (the neighborhood of the solution). The entropy of the initial region decreases linearly with the training sample size (hence the free energy $F_{0}=E_{0}-\theta S_{0}$ increases linearly). The free energy $F_{1}$ of the transition region changes slightly with the training sample size. Consequently, the grokking time decreases exponentially when the training sample increases (this agrees with observations in [9]).
Remark: Transition State Theory in catalysis and Grokking
Catalysis in chemistry works by increasing entropy $S_{1}$ and hence lowering free energy $F_{1}=E_{1}-\theta S_{1}$ of the transition state in Eyring’s formula $e^{-\beta(F_{1}-F_{0})}$ of the reaction rate.
For grokking — when the training sample increases, entropy of the initial state $S_{0}$ decreases and free energy $F_{0}=E_{0}-\theta S_{0}$ increases.
Both catalysis and grokking depend on the free energy difference $F_{1}-F_{0}$.
Therefore, increasing of the sample size in grokking works as catalysis in entropy driven chemical reaction.
@@ 4 Summary
We also have explained the grokking (delayed generalization) phenomenon and some properties of grokking observed in [9] as a manifestation of thermodynamic and kinetic phenomena — transition to generalization in grokking proceeds by entropic mechanism of chemical reaction, properties of grokking are explained by Brownian motion and Eyring’s formula. Delay of generalization in grokking is a result of transition through the barrier of entropy, and the delay is large when entropy of the initial state and therefore the barrier of free energy in Eyring’s formula is large.